OpenStudy (anonymous):

I am having such a hard time w/ this question, anyone? Helper will get a medal! Algebra: Option 1 (amount in $) year 1 = 1300 year = 2 1600 year 3 = 1900 Option 2 ($) year 1 = 1120 year 2 = 1254.40 year 3 = 1404.93 Part A: What type of function, linear or exponential, can be used to describe the value of the investment after a fixed number of years using option 1 and option 2? Explain your answer.

OpenStudy (anonymous):

Please ANYONE HELP ME!!!!! PLEASE!!!!!!

OpenStudy (anonymous):

@pgpilot326

OpenStudy (anonymous):

@AllTehMaffs please help me, I'm about to fail this test!

OpenStudy (anonymous):

Well it keeps going up by 300 in the 1st one. So that one has to be linear.

OpenStudy (anonymous):

Help, please? I'm really trying but I don't understand.

OpenStudy (anonymous):

Okay I can try. So it would be option one is linear, option 2 is exponential. Because linear goes up by an even number. 300 is even if you can get to it by an even number.

OpenStudy (anonymous):

Linear more specifically means that for the function that descirbes option 1, it can be expressed as \[f(t) = 1000 + 300t\] where t is in years

OpenStudy (anonymous):

Ok, and Option 2 is um, .....I don't know. ;__;

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

That one is exponential, trying to come up with its equation.

OpenStudy (anonymous):

So, is the question asking me which one is linear and which one is exponential?

OpenStudy (anonymous):

yeah. I thought that you had to come up with their functions but apparently that's not true.

OpenStudy (anonymous):

Oh. So, Option 1. is linear because the it is going evenly through out the 3 years, and Option 2. is exponential because it isn't going by a linear pattern.

OpenStudy (anonymous):

Would you mind helping me w/ Part B.? I won't bother you again!

OpenStudy (anonymous):

Yeah, I believe that suffices. And I'll try ^_^

OpenStudy (anonymous):

Awesome, thank you! Okay, here it is: Write one function for each option to describe the value of the investment f(n), in dollars, after n years. Option 1. I already know is f(x) = 1,000x +300. Could you help me find the function for Option 2?

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (anonymous):

I'm trying, but my brain has apparently shut off. First off, your first function is incorrect. It's \[f(t)=1,000 + 300t\] or, as you wrote it \[f(x) = 1,000 + 300x\] where year 1 is assigned to t=1

OpenStudy (anonymous):

it's increasing by 300 each time increment, so it's the 300 term that has the time variable attached to it.

OpenStudy (anonymous):

Oh, whoops. Here f(t) = 1,000 + 300t

OpenStudy (anonymous):

I understand, just made an error.

OpenStudy (anonymous):

Now, what next oh great one? :)

OpenStudy (anonymous):

Are you solving it? :3

OpenStudy (anonymous):

Trying, my brain is not cooperating with me today.

OpenStudy (anonymous):

Me too. Hmm...maybe I could help a little.

OpenStudy (anonymous):

@mebmod any suggestions?

OpenStudy (anonymous):

trying

OpenStudy (anonymous):

yay, okay, so it fits the form \[N = N_0(k^t)\]

OpenStudy (anonymous):

Awesome!!.....uh....hm....I think I get it, but would you mind explaining it to me, just to make sure?

OpenStudy (anonymous):

where k is some growth constant, and t is the time. For t=0 (year 1), we have our original population, but for t=1 (year 2) \[1254.4 = 1120(k^1)\] or just \[1254.4 = 1120\times k\] so what's k?

OpenStudy (anonymous):

.......I don't know. So sorry!

OpenStudy (anonymous):

WAIT

OpenStudy (anonymous):

HOLD ON A SECOND, I THINK I AM UNDERSTANDING.

OpenStudy (anonymous):

1.12?

OpenStudy (anonymous):

So if you were given \[1254.4=1120x\] how would you find x?

OpenStudy (anonymous):

I would divide both sides by 1,120.

OpenStudy (anonymous):

And then what? :)

OpenStudy (anonymous):

Then you just found k! ^_^ it's the same process 1254.4 = 1120k \[k=\frac{1254.4}{1120} = ?\]

OpenStudy (anonymous):

k = 1.12! :D

OpenStudy (anonymous):

So, a review: Option 1.)'s function would be f(t) = 1000 +300t. But here's the thing, it says it needs to be f(n) and then "n". What should I do?

OpenStudy (anonymous):

yup yup ^_^ So the exponential function for any time, t is \[f(t) = 1120k^t\] you can check that it gives the correct output for t=2 (year3) So first, I'm gonna make a correction, and say that at year 1, t=0, so the function turns into \[f(n) = 1300 + 300n\] All you're doing is adding 300 every year>1 so them saying f(n) is just a function where the thing that's going up is called n (maybe for numberof years). there wasn't some defined value for n, was there?

OpenStudy (anonymous):

n = years, and f(n) was dollars.

OpenStudy (anonymous):

Then yeah \[f(n) = $1200 + ($300/year)n\] where n is years.

OpenStudy (anonymous):

oops, sorry, should be 1300 for the first term

OpenStudy (anonymous):

Ok, so Option 1. is f(n) = 1300 +300n. Option 2.) is 1,254.40 = 1,120(n)

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Wait, when I do the math for the first options function, it gives me 1,900 for the second year. :(

OpenStudy (anonymous):

For option 1, remember that n=0 means year 1, so for years 2, n=1 Also, not quite right for option 2 Remember that it's exponential, so the n has to be in the exponent! ^_^ It's also still a function of n, so 1254.4 is just an output of the function. \[f(n) = 1120\times(1.12)^n\] where, again, n=0 corresponds to year 1, n=1 is year 2 n=2 is year 3 etc.

OpenStudy (anonymous):

I understand now! Awesome!

OpenStudy (anonymous):

1 more question, what would be the value of the investment f(n) for each options function? Dumb, I know.

OpenStudy (anonymous):

I don't understand the question. I'm bad with financial terms :/

OpenStudy (anonymous):

value is like the gains on the original investment, right?

OpenStudy (anonymous):

I don't know my friend, I don't know.

OpenStudy (anonymous):

Then your guess is as good as mine ^_^ Sorry :P

OpenStudy (anonymous):

It's cool, thank you so much for all the help!

OpenStudy (anonymous):

Very welcome. Glad that I could help! ^_^